A hot air balloon has a volume of 2000 m3 when fully inflated and the air inside the balloon is at atmospheric pressure (101,000 Pa). What is the mass of hot air inside the balloon if its temperature is 120°C? (Assume a molecular weight of 28.8 g/mole for air, which is about 80% N2 and 20% O2.)
Answer #1

We can solve this using Ideal Gas Equation and by finding the number of moles of air.

Ideal Gas Equation, PV = nRT

(101000)(2000) = n*8.314*393 {Temperature is in Kelvin, i.e. T = 273 + 120 = 393}

n = 61822.82 mol

Mass of air, M = n*(molecular weight)

M = 61822.82*28.8

M = 1780497.16 g = 1780.5 Kg

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.